Which of the following can be the sides of a right triangle?
2.5 cm, 6.5 cm, 6 cm.
In the case of right-angled triangles, identify the right angles

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let a = 2.5 cm, b = 6.5 cm, c = 6 cm
Let us assume the largest value is the hypotenuse side i.e. b = 6.5 cm.
Then, by Pythagoras theorem,
b^2=a^2+c^2
6.5^2= 2.5^2+ 6^2
42.25 = 6.25 + 36 42.25
= 42.25
The sum of square of two side of triangle is equal to the square of third side,
∴The given triangle is right-angled triangle.
Right angle lies on the opposite of the greater side 6.5 cm.

Thanks